OpenStudy (anonymous):

If the sin 90° is 1, then the cos ____ °= _____. 90° ; 1 over 2 0° ; square root 2 over 2 90° ; square root 3 over 2 0° ; 1

OpenStudy (anonymous):

Rule of trigonometric functions:- sin a = cos(90 - a) Here a = 90°. sin 90 = cos (90 - 90) 1 = cos 0. cos 0 = 1.

OpenStudy (anonymous):

Do you understand it now?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Okay well let me explain it to you.

OpenStudy (anonymous):

So is cos 0° = 1, because the angles are complementary